OpenStudy (anonymous):

g(t) = sqrt(2-t). Find g^-1(s)

OpenStudy (anonymous):

y = sqrt(x - 2) x = sqrt (y - 2) x^2 = y - 2 x^2 + 2 = y f^-1(x) = x^2 + 2 is this correct?

OpenStudy (anonymous):

I changed it from 2 -x to x -2 because I thought it would make it easier to solve

OpenStudy (anonymous):

yes